2. Show that in an open economy private saving has to be allocated between investment, government debt and purchases of wealth from foreigners.

That is, show that S = I + CA + (G T
Answer #1

As per National Income Identity:

Y = (C - T) + I + G + X - M

Y = C - T + I + G + X - M

(Y - C) = I + (X - M) + (G - T)

S = I + CA + (G - T), since CA = X - M